In what ratio should a 20% methyl alcohol solution be mixed with a 50% methyl alcohol solution so that the resultant solution has 40% methyl alcohol in it?
Answer by richwmiller(17219)

you are mixing two solutions to get a third.
so a+b=c
we don't know how much to mix nor how much we will end up with.
so we call the amount to mix x and y and the total will of course be x+y
.2x+.5y=.4(x+y)
2x+5y=4(x+y)
2x+5y=4x+4y
2x+y=4x
y=2x
so you will mix 2x for every y
y/x=2
x/y=1/2